OpenStudy (anonymous):

Physics - Medal+Fan A bowling ball is dropped from the top of a tower and takes 5.6 seconds to hit the ground. What is the displacement?

OpenStudy (campbell_st):

isn't displacement \[s = ut + \frac{1}{2} at^2\] u = initial velocity so u = 0 t = time and a is the acceleration, in this case due to gravity... so you should have a number for that and t = 5.6 just substitute and calculate

OpenStudy (anonymous):

I am new and trying to learn this how do I substitute and calculate

OpenStudy (campbell_st):

well \[s = 0 \times 5.6 + \frac{1}{2} \times a \times 5.6^2\] you need to substitute your value for a... acceleration due to gravity